Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

Unfair coin flip

Expert replies
by XLogic » Sun Apr 15, 2012 3:03 am
A bag contains 2 fair coins and 2 unfair coins. Both unfair coins have a 60 percent chance of showing up Heads if flipped. If I pick a coin from the bag and flip it 4 times, what is the probability that I will come up with tails 2 out of the 4 flips?
my post helped --> thank me!
don't thank me --> my post = what the..??
Join the discussion
Source: — Problem Solving |

by killer1387 » Sun Apr 15, 2012 3:22 am
XLogic wrote:A bag contains 2 fair coins and 2 unfair coins. Both unfair coins have a 60 percent chance of showing up Heads if flipped. If I pick a coin from the bag and flip it 4 times, what is the probability that I will come up with tails 2 out of the 4 flips?
probability of picking a fair coin= 2/4=1/2= probability of picking unfair coin

suppose the coin picked is fair one;

(1/2)*(1/2)^4= 0.5^5

coin picked is unfair one;

(1/2)*(0.6)^2*(0.4)^2

required probability = 0.03125+0.0288= 0.06005
Join the discussion

by Neo Anderson » Sun Apr 15, 2012 5:18 am
killer1387 wrote:
XLogic wrote:A bag contains 2 fair coins and 2 unfair coins. Both unfair coins have a 60 percent chance of showing up Heads if flipped. If I pick a coin from the bag and flip it 4 times, what is the probability that I will come up with tails 2 out of the 4 flips?
probability of picking a fair coin= 2/4=1/2= probability of picking unfair coin

suppose the coin picked is fair one;

(1/2)*(1/2)^4= 0.5^5

coin picked is unfair one;

(1/2)*(0.6)^2*(0.4)^2

required probability = 0.03125+0.0288= 0.06005
I did exactly the same thing whilst calculating the probability, except for one thing:-
In either case(fair of unfair coin), two tails can turn up in 4 flips in exactly 6 ways (HHTT, HTHT, TTHH, THTH, THHT and HTTH) so the total probability will be 6 times the above value!!

what say??
Join the discussion

by killer1387 » Sun Apr 15, 2012 7:29 am
Neo Anderson wrote:
killer1387 wrote:
XLogic wrote:A bag contains 2 fair coins and 2 unfair coins. Both unfair coins have a 60 percent chance of showing up Heads if flipped. If I pick a coin from the bag and flip it 4 times, what is the probability that I will come up with tails 2 out of the 4 flips?
probability of picking a fair coin= 2/4=1/2= probability of picking unfair coin

suppose the coin picked is fair one;

(1/2)*(1/2)^4= 0.5^5

coin picked is unfair one;

(1/2)*(0.6)^2*(0.4)^2

required probability = 0.03125+0.0288= 0.06005
I did exactly the same thing whilst calculating the probability, except for one thing:-
In either case(fair of unfair coin), two tails can turn up in 4 flips in exactly 6 ways (HHTT, HTHT, TTHH, THTH, THHT and HTTH) so the total probability will be 6 times the above value!!

what say??
i thought about it but cant reach a consensus logically so waiting for some comment describing the issue at hand.
Join the discussion

by shubham_k » Sun Apr 15, 2012 9:53 am
I think Xlogic is correct , the result has to be multiplied by 6 as they are six favorable sets. Killer1387 has obtained the answer without considering the number of ways it can be achieved in 4 tosses. or in other words he has got the result for a single set like getting only a HHTT combination. So the final answer is definitely 0.06005*6

Please take a moment to thank if you feel this post was helpful
Join the discussion

by XLogic » Sun Apr 15, 2012 1:16 pm
I did similar to Neo.

P(fair) = (1/2)^4 * 4C2
= (1/2)^4 * 4!/2!2!
= (1/2)^4 * 6

P(unfair) = (2/5)^2 * (3/5)^2 * 6
= (6/25)^2 * 6

P(unfair or fair) = (6/25)^2 * 6 + (1/2)^4 *6
= 6*[(6/25)^2 + (1/2)^4)]
my post helped --> thank me!
don't thank me --> my post = what the..??
Join the discussion

by killer1387 » Sun Apr 15, 2012 5:40 pm
I dont see as to why we need to consider arrangements here , we just have 2 tails out of 4.

btw whats the OA??
Join the discussion

by XLogic » Sun Apr 15, 2012 6:30 pm
@Killer, you don't just have 2 tails out of 4. You have X number of ways of getting 2 out of 4.

If we simplified the question to: If I tossed a fair coin 3 times, what is the probability of heads showing up exactly twice?

The question is how many ways is this possible:
HHT or
HTH or
THH

Therefore we get: (1/2 * 1/2 * 1/2) 3 times
3*(1/2)^3

If the question was: A fair coin is tossed three times, what is the probability that only the first 2 tosses end up with heads.

Then this is more restrictive: HHT = 1/2 * 1/2 * 1/2 = (1/2)^3

Thoughts??

OA: None
my post helped --> thank me!
don't thank me --> my post = what the..??
Join the discussion

by Neo Anderson » Mon Apr 16, 2012 12:22 am
XLogic wrote:@Killer, you don't just have 2 tails out of 4. You have X number of ways of getting 2 out of 4.

If we simplified the question to: If I tossed a fair coin 3 times, what is the probability of heads showing up exactly twice?

The question is how many ways is this possible:
HHT or
HTH or
THH

Therefore we get: (1/2 * 1/2 * 1/2) 3 times
3*(1/2)^3

If the question was: A fair coin is tossed three times, what is the probability that only the first 2 tosses end up with heads.

Then this is more restrictive: HHT = 1/2 * 1/2 * 1/2 = (1/2)^3

Thoughts??

OA: None
You have beautifully put it into words! Great Going!
Join the discussion